LASIK eye surgery uses pulses of laser light to shave off tissue from the cornea, reshaping it. A typical LASIK laser emits a 1.0-mm-diameter laser beam with a wavelength of 193 nm. Each laser pulse lasts 20ns and contains 1.7mJ of light energy.

What is the power of one laser pulse?

During the very brief time of the pulse, what is the intensity of the light wave?
Answer #1

2*10^-3/(19*10^-9) = 105kW
Intensity = 105/(0.001^2) = 105GW/m^2